Data shift report


Subject: Data shift report
From: ORSINI MASSIMO (Massimo.Orsini@lngs.infn.it)
Date: Fri Mar 03 2000 - 11:34:02 EST


Data shift report (Since 28-feb-2000 to 03-mar-2000)

-------------------------------------------------------------------------
  (by Nicola Zaccheo & Massimo Orsini)
  From run 19163 to 19190
 -------------------------------------------------------------------------
 During the week the acquisition has been running rather smoothly although
 on friday uVax1 and uVax2 didn't have enough resources to perform the usual
 Emily HV check.
 We stopped the run 19191 and we booted both uvaxes solving the problem.
   
    WFD warnings:

     run 19163 SM1CH6 --> 0-supp. off (evn.14)
     run 19167 SM2CH0 --> 78 NDATA -> all 0-supp. off
     run 19168 SM6CH34 --> big pulse from 6T02-0
     run 19169 SM2CH4 --> 0-supp. off (evn.139)
     run 19175 SM2CH? --> 0-supp. off (evn.16)
     run 19183 SM2CH0 --> 0-supp. off
     run 19184 SM2CH8 --> 0-supp. off
               SM6CH29 --> Ped. Var.
     run 19185 SM2CH0 --> 0-supp. off
     run 19187 SM2CH2 --> 0-supp. off
               SM2CH0 --> LED pulse
               SM2CH0 --> LED pulse
               SM4CH37 --> Big pulse
               SM4CH39 --> BIg pulse
     run 19188 SM2CH0 --> LED pulse
               SM2CH0 --> LED pulse
               SM2CH0 --> LED pulse
     run 19189 SM2CH4 --> 0-supp. off
     run 19190 SM1CH7 --> Big pulse
               SM2CH0 --> LED pulse
               SM2CH0 --> LED pulse

    TRIGGER RATE (hot tanks):
              ERP --> 3B07~8mHz, 6T07~9mHz
              LIP --> 2E10~12mHz, 3E04~6.5mHz, 6T12~11mHz
              TOHM --> 1T11~20mHz, 2T16~15mHz, 3T01~9mHz, 3T12~50mHz(!)
                       4T02~10mHz, 4T11~10mHz, 5T02~15mHz, 6B09~4mHz,
                       6B13~6mHz, 6T11~9mHz

  LAMOSSKA problems:
              only not justified LAM50 triggers as following:
              run 19165 --> 1 evn
              run 19167 --> 1 evn
              run 19168 --> 2 evn
              run 19169 --> 1 evn
              run 19170 --> 1 evn + 1 evn from SM4 (LAM41)
              run 19171 --> 1 evn
              run 19172 --> 1 evn
  LAMOSSKA triggers are alive from SM2 to SM6 and all of them gave the early
  stop to the WFD's but from run 19166 untill run 19188 we didn't get any
  LAMOSSKA triggers from SM1 side 0.

    CSPAM/ERP efficiency:
              3B07 ~ 40% but 6T07 is now @~83%(!)
                     sometimes 4E14 @50% and 3C01 @70%
        NOTE: 3B07 has not a good efficiency because it has the former
              body leaker.

    LIP/ERP efficiency:
               1B16 sometime has big ineff. (see run 19171-> 3iER:13, iLIP:1)
  
    SPU: still hits on ERPI and ERPII on uVax1. See, for instance, run 19175
         (ERPI:4, ERPII:4)
         Still LED hits on uVax1.

    STOP MASTER: there are no Q's from STOM on SM5 in every runs. They range
                 from 4 to more than 20. They seem to be present on one-for-
                 event base also during events in wich the STPM on SM5 shoud
                 not be read(!)
                 Yiannis and me checked some of these events w/o finding
                 simply explainations..
                 We are getting some fake hits (~30 per run) from
                 SM4%1@CITm and ~10 fake hits from SM4%3@WEST.
                  
     CALIBRATIONS performed on uVAX3:
                   
             LED/TDC, LASER LOWER, LASER ATTICO and FMT.
   
     HARDWARE INTERVENTIONS:
                    
   We had to turn off both VXMACA/B in order to allow the UPS maintenance.
   For this reason we had to connect the ethernet bridges and all DEC servers
   to the BLINDO SBARRA of Macro using some power cord extenders, running
   the main acquisition on VSMAC1 after having plugged it to the Macro main
   power line.
   WSGS04 (PISA monitor) also needed to be turned off and booted in order to
   have the supernova WATCH alive.
   During those operations VSMAC3 monitor broke down and our run coordinator
   is now trying to get a new one.
   The LNGS technicians replaced also two PILLER batteries but they didn't
   find out any UPS problems.

  MACRO DATA REPROCESSING: (by Roberto Giuliani)
   
    DD093,94 are now still in progress.
    From DD077 to DD092 all ZEBRA files were succesfully reprocessed.
    A new TA90 directory (DDM14) was already copied/backupped on tapes.
 
                        -------- THE END -------



This archive was generated by hypermail 2a24 : Fri Mar 03 2000 - 11:34:08 EST